Price per bushelQuantity Demanded (bushels)Quantity Supplied (bushels)
US$240,0000
436,0004,000
630,0008,000
824,00016,000
1020,00020,000
1218,00028,000
1412,00036,000
166,00040,000


2. Refer to Table 1-2. The table contains information about the wheat market. Use the table to answer the following questions.


a. What are the equilibrium price and quantity of wheat?


b. Suppose the prevailing price is US$6 per bushel. Is there a shortage or a surplus in the market?


c. What is the quantity of the shortage or surplus?


d. How many bushels will be sold if the market price is US$6 per bushel?


e. If the market price is US$6 per bushel, what must happen to restore equilibrium in the market?


f. At what price will suppliers be able to sell 36,000 bushels of wheat?


g. Suppose the market price is US$14 per bushel. Is there a shortage or a surplus in the market?


h. What is the quantity of the shortage or surplus?


i. How many bushels will be sold if the market price is US$14 per bushel?


j. If the market price is US$14 per bushel, what must happen to restore equilibrium in the market?




Table 1-2 Quantity Quantity Demanded supplied Price per bushel (bushels) bushels) US$2 40,000 10 36,000 4,000 4 6 8,000 30,000 8 24,000 16,000 10 20,000 20,000 18,000 28,000 12 14 12,000 36,000 6,000 16 40,000
Answer #1

a) Equilibrium is where QD=QS

Equilibrium price = 10

Equilibrium quantity = 20000

b) At P =6

QD=30000

QS=8000

So, there is a shortage

c) Shortage = 30000-8000 = 22000

d) Bushels sold = quantity supplied = 8000

e) The price must rise so that the quantity supplied increases

f) At P = 4

g) When P =14, QD = 12000, QS = 36000

So, there is a surplus in the market

h) Surplus = 36000-12000 = 24000

i) Quantity sold = quantity demanded = 12000

j) The price must fall so that the demand increases